OpenStudy (anonymous):

Solve 3x^2 + 15 = -6x

OpenStudy (anonymous):

Answer choices:

OpenStudy (anonymous):

@Jaylikescookies @phi @petiteme @Destinymasha @Euler271 @Hailey553 Help please :)

OpenStudy (anonymous):

These involve imaginary numbers.. I suck at imaginary numbers. Sorry :c

OpenStudy (anonymous):

@Jaylikescookies Lol that's okay, I suck at math in general

OpenStudy (anonymous):

\[3x^2 + 15 = -6x\] start with \[3x^2+6x=-15\] then divide by 3 to get \[x^2+2x=-5\]

OpenStudy (anonymous):

probably easiest to complete the square \[(x+1)^2=-5+1=-4\] so \[x+1=\pm\sqrt{-4}=\pm2i\]

OpenStudy (anonymous):

subtract 1 and get \[x=-1\pm2i\]

OpenStudy (anonymous):

Thank you! ..but that's that an answer choice..

OpenStudy (anonymous):

i am sticking with my answer

OpenStudy (anonymous):

in fact it comes with a money back guarantee

OpenStudy (anonymous):

Lol but it's not an answer choice! What do I do?

OpenStudy (anonymous):

i am only authorized to give answers and help with math problems

OpenStudy (anonymous):

I know, but..I think you're wrong :P These are the answer choices:

OpenStudy (anonymous):

@satellite73